OpenStudy (anonymous):

Solve the equation. log(x + 9) – log x = 3

OpenStudy (campbell_st):

ok... do you know about the log law for division...?

OpenStudy (campbell_st):

\[\log (\frac{a}{b} )= \log(a) - \log(b) \]

OpenStudy (campbell_st):

you also need to know about raising the log to a power of the base \[a^{\log_{a}(x)} = x\] and you appear to be working with base 10 logs so you have \[\log(\frac{x + 9}{x} )= 3\] raise them to the power of 10 so \[\frac{x + 9}{x} = 10^3\] now you can solve for x